# Generative AI for Predicting 2D and 3D Wildfire Spread: Beyond Physics-Based Models and Traditional Deep Learning

### v1

@@ 3 Emerging Generative AI Models and Their Advantages
Since around 2018 and especially in the early 2020s, generative AI models have experienced rapid advancements and widespread adoption across diverse fields, marking a significant shift in the landscape of artificial intelligence (Brown et al., 2020). Broadly defined, generative AI refers to a class of machine learning models that are capable of learning the underlying distribution of data and generating new, realistic content—such as text, images, audio, or even simulations—that resembles the training data (Goodfellow et al., 2014; Kingma et al., 2013; Ho et al., 2020). These models and their architectures have been increasingly experimented with and adopted to solve complex problems across various scientific disciplines.
@@ 3.1 Generative AI Applications in Environmental and Urban Sciences
The recent emerging generative AI models go beyond traditional predictive analytics by creating novel outputs, making them highly valuable in data-scarce environments or in domains requiring scenario generation, synthesis, or simulation. Generative AI encompasses several prominent categories, including GANs, VAEs, Autoregressive Models (AR), Diffusion Models, and Flow-based Models (Rezende & Mohamed, 2015; Vaswani et al., 2017; Kingma & Dhariwal, 2018). Within these categories, more specialized types have emerged: GANs consist of a generator and discriminator in adversarial training to synthesize realistic data (Goodfellow et al., 2014); VAEs combine probabilistic encoding and decoding for efficient latent space learning (Kingma et al., 2013); autoregressive models like GPT-4 predict future tokens in sequences for high-quality text generation (Brown et al., 2020); diffusion models such as Stable Diffusion and denoising diffusion probabilistic models (DDPMs) generate content by gradually denoising random noise into coherent samples (Ho et al., 2020; Rombach et al., 2022); and flow-based models leverage invertible transformations for exact likelihood estimation and sample generation (Dinh et al., 2017).
@@
Each generative AI model and its architecture offers distinct mathematical properties and advantages depending on the application context. The rise of these models has revolutionized data science and AI by enabling machines to not only interpret and predict but also to imagine and create, thereby unlocking new opportunities across fields including hazard prediction, urban planning, climate science, decision support, design automation, and digital twin construction (Xu et al., 2024c; Ma et al., 2024). In the environmental hazard domain, Ma et al. (2024) presents a comprehensive review of how generative AI addresses longstanding challenges in data availability, quality, and resolution. These models learn high-dimensional probability distributions from limited samples, making them ideal for data-scarce applications such as geohazards, hydrometeorology, and climate-related analysis. By generating physically consistent synthetic data—ranging from downscaled meteorological fields to simulated landslide and seismic imagery—GenAI enhances forecasting accuracy, susceptibility mapping, and early warning systems, driving more reliable and scalable hazard modeling frameworks. In the urban management sector, Xu et al. (2024c) conducts a scoping review on integrating GenAI into urban digital twins, highlighting its transformative role in automating the generation of high-quality urban data, hypothetical planning scenarios, and 3D city models. These capabilities help overcome major challenges related to data sparsity, simulation scalability, and design complexity. Through the synthesis of multi-modal urban data and simulation of complex dynamics, GenAI enhances real-time decision support, predictive analytics, and participatory planning across sectors such as transportation, energy, water, and infrastructure. The convergence of GenAI with digital twin technologies represents a paradigm shift toward intelligent, adaptive, and inclusive urban solutions. Furthermore, growing research extends GenAI applications into other urban subsystems, including logistics optimization, intelligent transportation systems, and domain-specific knowledge generation (Tupayachi et al., 2024; Xu et al., 2024e; Taiwo et al., 2025).
3.2 Advantage over Traditional Deep Learning Models
Building on the limitations identified in the existing wildfire prediction literature, we conducted a comparative analysis of traditional deep learning models and generative AI approaches. As summarized in Section 2.3, our findings underscore the key advantages of generative AI models in addressing the constraints of conventional methods. These advantages are outlined in the list below, and a detailed comparison is provided in Table 3, which maps each limitation (L1–L6) discussed in Section 2.3 to corresponding strengths of generative AI.
@@ 3.2 Advantage over Traditional Deep Learning Models
In contrast to many traditional deep learning models, generative AI models such as VAEs and diffusion models produce probabilistic outputs that inherently capture uncertainty in predictions—an essential capability for high-risk applications like wildfire forecasting (Kingma et al., 2013; Ho et al., 2020). This feature can be leveraged to address the limitation L1.
@@
Transformer-based architectures outperform recurrent neural networks (RNNs) in modeling long-range dependencies in sequential data, making them well-suited for tracking wildfire dynamics over extended time periods (Vaswani et al., 2017; Brown et al., 2020) The capability could be harnessed to tackle the limitation L2.
@@
Generative AI models, particularly those based on transformers and diffusion techniques, excel at integrating heterogeneous data sources (e.g., satellite imagery, meteorological variables, and point clouds), enabling more robust 2D and 3D wildfire forecasting through a unified framework (Rombach et al., 2022; Radford et al., 2021) to address the limitation L3.
@@
Generative AI models such as GANs and diffusion models can produce synthetic yet realistic wildfire progression data, providing valuable training samples in data-scarce scenarios, facilitating missing data imputation to enhance data quality, and supporting the simulation of extreme or hypothetical conditions (Goodfellow et al., 2014; Dhariwal and Nichol, 2021). This capability can be leveraged to address Limitations L4 and L5, as well as to enable scenario generation for simulating hypothetical wildfire events.
@@
Many generative AI models, such as VAEs and GANs, rely on latent spaces and latent vectors to operate, where complex data relationships are encoded in lower-dimensional representations (Kingma et al., 2013; Goodfellow et al., 2014). These latent variables can be visualized and analyzed to enhance the explainability, interpretability, and trustworthiness of the models in tasks such as classification, prediction, clustering, and data generation (Xu et al., 2024a), thereby addressing Limitation L6.
@@
4 Review Strategy and Methodologies
This study employs a human–AI collaborative approach by leveraging an automated pipeline powered by a LLM to systematically retrieve relevant research articles from multiple academic databases (Xu et al., 2024b). Guided by user-defined search queries, the system automates literature retrieval, characterization, and knowledge synthesis to generate bibliometric summaries. These summaries provide insights into how various machine learning algorithms, and more recently, generative AI models, have been applied to support fire spread prediction and propagation modeling, often in conjunction with complementary computing and informatics technologies.
To conduct a critical review of deep learning and generative AI , we developed a series of targeted search queries to perform a comprehensive literature review across both the IEEE Xplore and Scopus databases. Our search strategy was designed to address two primary objectives: first, to conduct a broad initial search that captures a wide spectrum of studies applying artificial intelligence and deep learning techniques to bushfire spread and propagation; and second, to narrow the focus by identifying recent and emerging research that specifically explores the use of generative AI models—such as VAEs, GANs, Transformers, and diffusion models—for wildfire prediction. The detailed search queries used for each phase of the review are presented in Figure 1.
4.1 LLM-Powered Literature Characterization
In addition to the identified articles, we utilized an LLM-powered tool developed in our previous work (Xu et al., 2024b) to characterize the literature by extracting taxonomies from each paper’s abstract using a combined process of Scientific Discourse Tagging (SDT) and Named Entity Recognition (NER). These extracted taxonomies were then classified based on an existing body of knowledge. A bibliometric summary is presented in Figure 2, which visualizes the number of publications employing various machine learning and deep learning models. The models are grouped into three categories: traditional machine learning, deep learning, and generative AI, and are displayed as bar charts. Figure 3 presents a more detailed characterization and summary of the literature, mapping the number of publications that apply various categories and types of machine learning and deep learning models across different wildfire prediction application areas over the years. The classification of the literature is conducted using Sentence Transformers, which align the extracted taxonomy from each article with the existing body of knowledge (Xu et al., 2024b). This alignment is performed using a cosine similarity threshold that compares the literature-derived taxonomy with named entities and definitions defined in a domain ontology, which was constructed based on comprehensive literature reviews and authoritative textbooks. In this study, a cosine similarity threshold of 0.7 is used to ensure reliable knowledge alignment, allowing literature to be classified under a specific methodology or application area only when its extracted taxonomy is highly semantically relevant to a corresponding definition in the existing body of knowledge.
@@ 4.1 LLM-Powered Literature Characterization
Figure 2: Distribution of literature counts across various machine learning and deep learning model types, categorized by application areas. The analysis was conducted using an LLM-powered research tool.
Figure 3: Classification of existing deep learning literature in wildfire science by model type, application area, and publication year, generated using an LLM-powered research tool.
Based on our bibliometric summary of articles retrieved from IEEE Xplore and Scopus using a general generative AI query, we find that applications of generative AI in wildfire prediction remain limited, primarily focusing on GANs and transformers, as shown in Figure 3. Although other generative models could similarly analyze or synthesize relevant environmental data, this gap highlights significant opportunities for future research in the field.
5 State-of-the-Art Generative AI Applications in Wildfire Management
This section provides an in-depth review of current applications of generative AI in the wildfire prediction domain. The literature discussed is drawn from a refined subset of 17 research articles identified through IEEE Xplore and Scopus queries, as illustrated in Figure 1. The refinement process focuses on selecting studies that employ generative AI models as the core research methodology for wildfire-related analysis, rather than those using generative models solely for generic data processing or preprocessing tasks. Beyond the LLM-powered literature characterization, we also conducted a manual review and validation to filter out 1 research articles that claimed to use generative AI based on general rationale, but did not actually implement any generative AI models or architectures. Another 3 articles were founding using the diffusion models from the environmental science sector, rather than using the stable diffuser, which is one type of generative AI models. In addition, we also manually examined and removed 1 article from the initial selection that is retracted by the journal.
Through the following subsections, we conduct a structured and in-depth review on 11 existing studies that have explored the use of generative AI models for wildfire management.

### v2

@@ 3 Emerging Generative AI Models and Their Advantages
Since around 2018 and especially in the early 2020s, generative AI models have experienced rapid advancements and widespread adoption across diverse fields, marking a significant shift in the landscape of artificial intelligence Brown et al. (2020). Broadly defined, generative AI refers to a class of machine learning models that are capable of learning the underlying distribution of data and generating new, realistic content—such as text, images, audio, or even simulations—that resembles the training data Goodfellow et al. (2014); Kingma et al. (2013); Ho et al. (2020). These models and their architectures have been increasingly experimented with and adopted to solve complex problems across various scientific disciplines.
@@ 3.1 Generative AI Applications in Environmental and Urban Sciences
The recent emerging generative AI models go beyond traditional predictive analytics by creating novel outputs, making them highly valuable in data-scarce environments or in domains requiring scenario generation, synthesis, or simulation. Generative AI encompasses several prominent categories, including GANs, VAEs, Autoregressive Models (AR), Diffusion Models, and Flow-based Models Vaswani et al. (2017); Brown et al. (2020); Kingma et al. (2013). Within these categories, more specialized types have emerged: GANs consist of a generator and discriminator in adversarial training to synthesize realistic data Goodfellow et al. (2014); VAEs combine probabilistic encoding and decoding for efficient latent space learning Kingma et al. (2013); autoregressive models like GPT-4 predict future tokens in sequences for high-quality text generation Brown et al. (2020); diffusion models such as Stable Diffusion and denoising diffusion probabilistic models (DDPMs) generate content by gradually denoising random noise into coherent samples Ho et al. (2020); Rombach et al. (2022); and flow-based models leverage invertible transformations for exact likelihood estimation and sample generation Dinh et al. (2017).
@@
Each generative AI model and its architecture offers distinct mathematical properties and advantages depending on the application context. The rise of these models has revolutionized data science and AI by enabling machines to not only interpret and predict but also to imagine and create, thereby unlocking new opportunities across fields including hazard prediction, urban planning, climate science, decision support, design automation, and digital twin construction Xu et al. (2024c); Ma et al. (2024). In the environmental hazard domain, Ma et al. (2024) presents a comprehensive review of how generative AI addresses longstanding challenges in data availability, quality, and resolution. These models learn high-dimensional probability distributions from limited samples, making them ideal for data-scarce applications such as geohazards, hydrometeorology, and climate-related analysis. By generating physically consistent synthetic data—ranging from downscaled meteorological fields to simulated landslide and seismic imagery—GenAI enhances forecasting accuracy, susceptibility mapping, and early warning systems, driving more reliable and scalable hazard modeling frameworks. In the urban management sector, Xu et al. (2024c) conducts a scoping review on integrating GenAI into urban digital twins, highlighting its transformative role in automating the generation of high-quality urban data, hypothetical planning scenarios, and 3D city models. These capabilities help overcome major challenges related to data sparsity, simulation scalability, and design complexity. Through the synthesis of multi-modal urban data and simulation of complex dynamics, GenAI enhances real-time decision support, predictive analytics, and participatory planning across sectors such as transportation, energy, water, and infrastructure. The convergence of GenAI with digital twin technologies represents a paradigm shift toward intelligent, adaptive, and inclusive urban solutions. Furthermore, growing research extends GenAI applications into other urban subsystems, including logistics optimization, intelligent transportation systems, and domain-specific knowledge generation Tupayachi et al. (2024); Xu et al. ; Taiwo et al. (2025).
3.2 Theoretical Foundations of Fire Spread Modeling for Generative AI Integration
Generative AI provides powerful computational tools to simulate, model, and predict natural phenomena, but it still relies on existing physical laws and fundamental principles. To explore how emerging generative AI architectures can enhance or redefine current wildfire spread simulations and their core processes, it is crucial to first revisit the fundamental principles underlying fire spread modeling. At their core, most wildfire spread simulators are built upon two foundational computational paradigms: Huygens’ principle-based wavefront expansion methods and grid-based local interaction models. Each embodies a distinct conceptual strategy to replicate the spatiotemporal dynamics of fire growth, offering complementary strengths and insights for different use cases Sun et al. (2023).
Huygens’ principle-based methods simulate wildfire spread by borrowing ideas from wave physics, focusing on how the physical process propagates over time. They treat every point along the fire edge as a new ignition point that pushes the fire outward, often forming an ellipse shaped by wind, slope, and fuel Anderson et al. (1982). This helps capture how fires stretch and move in different directions. Many widely used systems, like Prometheus, FARSITE, and FlamMap, rely on this approach to keep updating the fire boundary over time Tymstra et al. (2010); Mitsopoulos et al. (2014). On the other hand, grid-based models focus on the characterization and connectivity of individual spatial or volumetric units created by dividing the landscape into a grid of small squares (2D cells) or cubes (3D voxels). This grid provides a spatial and temporal framework over which physical processes, such as wildfire, can propagate Barton et al. (2020). Each unit records its combustion state—unburned, burning, or burned—and captures local fuel and environmental characteristics. Fire spreads by following local transition rules influenced by fuel type, wind, slope, and stochastic effects Sun et al. (2023). Tools like Cell2Fire and hybrid models that combine grid methods with physics or machine learning build on this idea Rui et al. (2018); Xu et al. (2022). While often exploratory, grid-based models are valuable for studying detailed fire patterns, testing new hypotheses, and extending simulations into 3D to better capture fire behavior.
From a computational standpoint, generating large-scale grid-based models often requires substantial resources, including high RAM and GPU memory. Initializing large, connected grids in parallel environments can also incur significant time costs due to communication and load-balancing overhead. Moreover, both Huygens’ principle-based and grid-based methods typically depend on integrating multiple layers of GIS and meteorological data—such as topography, fuel attributes, and weather conditions—to achieve accurate physical or empirical simulations. Consequently, simulating wildfire spread over large areas or extended timeframes becomes highly resource-intensive, significantly increasing inference time and limiting the practicality of these approaches for rapid or wide-area fire behavior forecasting.
To provide a clearer theoretical foundation regarding how these generative approaches can be integrated into the Huygens’ principle-based methods and the grid-based model, we first outline the core mathematical formulations that govern their learning objectives and elucidate their ability to capture both local and non-local fire spread dynamics. Conceptually, VAEs, diffusion models, and Transformers can be seamlessly integrated with these core fire spread simulation principles by operating over localized kernels of neighboring cells surrounding active burning fronts. This architecture enables the models to learn spatial and temporal transitions of adjacent unburned cells, effectively guided by the propagation dynamics derived from physics-based simulation outputs or real-world data. Simultaneously, by embedding fuel characteristics and environmental covariates within their latent representations or as embeddings, these models capture the complex, multifactorial behavior underlying wildfire propagation, thus offering a more computationally efficient and rapid means to model and predict fire spread beyond traditional physical rules Yang et al. (2024); Joshi (2025).
Conceptually, VAEs optimize a stochastic latent-variable model by maximizing the evidence lower bound (ELBO), which balances reconstruction accuracy against divergence from a prior distribution and thus enables learning complex distributions over fine-scale spatial fire spread patterns. In contrast, diffusion models learn to reverse a gradual noising process through a denoising objective, allowing them to generate plausible next-timestep fire perimeters defined by Huygens’ principle and conditioned on environmental factors Zhihan (2024). Meanwhile, transformers use self-attention mechanisms to dynamically capture long-range spatial and temporal dependencies, making them well-suited to learn how fires propagate across heterogeneous landscapes without explicitly encoding all physical interactions. In the following section, we examine the mathematical formulations of each generative AI architecture and explore, at a theoretical level, how they can be integrated with the core principles of fire spread prediction.
3.2.1 VAE
Diving into the details, VAEs optimize a stochastic latent-variable model by maximizing the ELBO, which is represented in the following equation that balances reconstruction fidelity with regularization toward a prior, enabling VAEs to have the potential ability learn distributions over plausible local fire spread patterns Kingma and Welling (2014).
$\mathcal{L}_{\text{VAE}}=\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x})}\!\big{[}\log p_{\theta}(\mathbf{x}|\mathbf{z})\big{]}-D_{\mathrm{KL}}\!\big{(}q_{\phi}(\mathbf{z}|\mathbf{x})\,\|\,p(\mathbf{z})\big{)},$ (1)
In the equation 3.2.1, $x$ represents the observed data—such as spatial snapshots of wildfire perimeters or voxel grids showing burned states—while $z$ denotes latent variables that capture hidden influences like wind, fuel, and terrain effects on fire spread. The encoder $q_{\phi}(z|x)$, parameterized by $\phi$, approximates the posterior distribution over these latent factors given the observed fire data, and the decoder $p_{\theta}(x|z)$, parameterized by $\theta$, reconstructs or generates fire spread patterns from them. The model optimizes the evidence lower bound (ELBO), balancing the reconstruction term $\mathbb{E}_{q_{\phi}(z|x)}[\log p_{\theta}(x|z)]$, which ensures the decoded spread closely matches the observed data, and the KL divergence term $D_{KL}(q_{\phi}(z|x)\|p(z))$, which regularizes the latent space toward a simple prior $p(z)$. In wildfire applications, this enables VAEs to learn meaningful compressed representations of local fire dynamics and to generate realistic alternative spread scenarios that remain consistent with underlying physical and environmental processes.
3.2.2 Diffusion Models
Diffusion models learn to reverse a gradual noising process, often through a simplified denoising objective (as shown in equation 3.2.2), allowing them to have the ability to generate realistic next-step fire perimeters conditioned on environmental features Ho et al. (2020).
$\mathcal{L}_{\text{DM}}=\mathbb{E}_{\mathbf{x},\epsilon,t}\!\!\left[\big{\|}\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},t)\big{\|}^{2}\right],$ (2)
In this equation, $x$ denotes the original data sample, which in wildfire modeling could represent a spatial map or voxel grid of a current fire perimeter or fuel state. The variable $t$ indexes the timestep within the diffusion process, tracking the progression of noise addition or removal. Correspondingly, $x_{t}$ is the noisy version of the data at step $t$, simulating partially corrupted fire perimeter states. The term $\epsilon$ represents the actual noise sampled and applied during the forward diffusion process, while $\epsilon_{\theta}(x_{t},t)$ is the model’s prediction of this noise, parameterized by $\theta$. By minimizing the difference between the true noise and the predicted noise across all samples, noise levels, and timesteps—captured by the expectation $\mathbb{E}_{x,\epsilon,t}$—the model learns to effectively reverse the noising process. In wildfire applications, this can potentially enable the generation of realistic next-timestamp fire perimeters conditioned on environmental features, by iteratively refining uncertain or noisy predictions into plausible spread scenarios that align with learned data distributions and underlying physical patterns.
3.2.3 Transformers
Finally, transformers use self-attention to model long-range spatial or temporal dependencies (as depicted in equation 3.2.3), which enables dynamic context aggregation across heterogeneous landscapes or multi-scale temporal sequences Vaswani et al. (2017).
$\text{Attention}(Q,K,V)=\text{softmax}\!\!\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V,$ (3)
In this equation, $Q$, $K$, and $V$ represent the query, key, and value matrices, respectively. These are learned projections of the input data that capture different spatial or temporal features. In wildfire spread modeling, for example, $Q$ might encode the characteristics of a focal location or time step—such as current fuel moisture, wind conditions, or ignition state—while $K$ and $V$ capture information from surrounding spatial cells or earlier temporal states. The operation $\mathrm{softmax}!\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)$ computes attention weights by measuring the similarity between the query and the keys, normalized by the dimension $d_{k}$ to ensure stable gradients. These weights then aggregate the values $V$, producing a context-aware representation that dynamically integrates long-range dependencies. In the context of fire spread, this mechanism could be incorporated into a grid-based framework to learn how a cell’s or voxel’s future state is influenced by conditions across heterogeneous landscapes and multi-step temporal interactions, effectively complementing classical Huygens’ principle and traditional grid-based simulators by automatically capturing complex, data-driven spatial and temporal relationships.
@@ 3.2.3 Transformers
3.3 Advantage over Traditional Deep Learning Models
Building on the limitations identified in the existing wildfire prediction literature, we conducted a comparative analysis of traditional deep learning models and generative AI approaches. As summarized in Section 2.3, our findings highlight the key advantages of generative AI models in overcoming the limitations of conventional methods. These advantages are outlined below, with a detailed comparison presented in Table 3, which maps each limitation (L1–L6) discussed in the previous section to the corresponding strengths of generative AI.
@@ 3.3 Advantage over Traditional Deep Learning Models
In contrast to many traditional deep learning models, generative AI models such as VAEs and diffusion models produce probabilistic outputs that inherently capture uncertainty in predictions—an essential capability for high-risk applications like wildfire forecasting Kingma et al. (2013); Ho et al. (2020). This feature can be leveraged to address the limitation L1.
@@
Transformer-based architectures outperform recurrent neural networks (RNNs) in modeling long-range dependencies in sequential data, making them well-suited for tracking wildfire dynamics over extended time periods Vaswani et al. (2017); Brown et al. (2020) The capability could be harnessed to tackle the limitation L2.
@@
Generative AI models, particularly those based on transformers and diffusion techniques, excel at integrating heterogeneous data sources (e.g., satellite imagery, meteorological variables, and point clouds), enabling more robust 2D and 3D wildfire forecasting through a unified framework Rombach et al. (2022); Radford et al. (2021) to address the limitation L3.
@@
Generative AI models such as GANs and diffusion models can produce synthetic yet realistic wildfire progression data, providing valuable training samples in data-scarce scenarios, facilitating missing data imputation to enhance data quality, and supporting the simulation of extreme or hypothetical conditions Goodfellow et al. (2014); Dhariwal and Nichol (2021). This capability can be leveraged to address Limitations L4 and L5, as well as to enable scenario generation for simulating hypothetical wildfire events.
@@
Many generative AI models, such as VAEs and GANs, rely on latent spaces and latent vectors to operate, where complex data relationships are encoded in lower-dimensional representations Kingma et al. (2013); Goodfellow et al. (2014). These latent variables can be visualized and analyzed to enhance the explainability, interpretability, and trustworthiness of the models in tasks such as classification, prediction, clustering, and data generation Xu et al. (2024a), thereby addressing Limitation L6.
@@
4 Review Strategy
To ensure methodological rigor and transparency, this study adopts a systematic review strategy combined with narrative synthesis. By systematically formulating search queries and applying them across well-established academic databases—including IEEE Xplore and Scopus—we identified and screened relevant literature in a structured, reproducible manner (as depicted in Figure 1). This approach enabled us to comprehensively map existing research on the application of artificial intelligence to wildfire spread prediction, while also allowing us to iteratively narrow our scope to focus on the emerging role of generative AI models. Following initial retrieval and screening, we employed a narrative synthesis to qualitatively analyze and compare selected studies, examining their methodological frameworks, data sources, and reported predictive performances. This process provided nuanced insights into how diverse deep learning and generative architectures have been leveraged to advance fire propagation modeling.
In addition to the identified articles, we employed an LLM-powered tool from our previous work Xu et al. (2024b) to characterize the literature by extracting taxonomies from each paper’s abstract using a combined process of Scientific Discourse Tagging (SDT) and Named Entity Recognition (NER). These taxonomies were then classified against an existing body of knowledge. Specifically, we used Sentence Transformers to align each article’s extracted taxonomy with domain knowledge by applying a cosine similarity threshold, comparing entities to definitions from a domain ontology built through comprehensive literature reviews and authoritative textbooks. A similarity threshold of 0.7 was chosen to ensure robust alignment, classifying papers under specific methodologies or application areas only when their extracted taxonomy showed high semantic relevance to existing definitions. For example, this process enabled classification of literature into application areas such as fire spread prediction, detection and monitoring, and risk assessment, as illustrated in Figure 1.
@@ 4 Review Strategy
5 Generative AI Applications in Wildfire Management
Through the following subsections, we present a structured, in-depth review of 11 existing studies that have explored the use of generative AI models for wildfire management. This review is organized according to the application areas shown in Figure 1, which were identified through keyword-based literature characterization. Our primary aim is to examine how generative AI has been applied to support fire prediction, while also highlighting related areas such as fire monitoring and risk mapping. We focus specifically on evaluating the prediction accuracy, inference time, and computational efficiency of these generative AI models compared to traditional methods used as benchmarks.